OpenStudy (perl):

The force of attraction on an object below earth's surface is directly proportional to it's distance from earth's center.Find the work done in moving a weight of w lb located A miles below earth's surface up to the surface itself. Assume earth's radius is a constant r mi.

OpenStudy (perl):

I would like someone to check this over. The work I have done so far: F = k*x Now the question is a little confusing, how to find k. If we assume that it is w lb at the surface of earth then w = k* r k = w/r So then we integrate \[\int\limits_{r-a}^{r}\frac{w}{r}xdx= \frac{w(2ar-a^2)}{2r} \]

OpenStudy (perl):

the book answer was \[\frac{w(2ar-a^2)}{2}\]

OpenStudy (perl):

@Michele_Laino maybe you could look at it

OpenStudy (anonymous):

w is GMm/r^2 so the extra r you are putting up there in integration is absurd

OpenStudy (michele_laino):

please note that force F is given by the subsequent fromula: F = k*r so the differential work dW done by gravity acting inside the earth, is : dW = Fdr dW = kr*dr where dr is the differential of r, so we have: \[W = \int\limits_{r - a}^a {k\,r\,dr} = \frac{k}{2}\left. {{r^2}} \right|_{r - a}^a = \frac{k}{2}\left( {2ar - {a^2}} \right)\] now, the gravity g(r) inside the earth is giveen by the subsequent formula: \[g\left( r \right) = {g_0}\frac{r}{a}\] where g_0 is the gravity onsurface of earth, so we can write: \[k \cdot r = m{g_0}\frac{r}{a} = w\frac{r}{a}\] then we get: \[k = \frac{w}{a}\] and substituting into the formula for the work done W, we can write: \[W = \frac{k}{2}\left( {2ar - {a^2}} \right) = \frac{k}{2}a\left( {2r - a} \right) = \frac{w}{2}\left( {2r - a} \right)\] Note that W<0 since W is the work done against the earth gravity.

OpenStudy (michele_laino):

please I have made an error, since W>0, not W<0

OpenStudy (perl):

@Michele_Laino i think there is a mistake , the integral should be from r-a to r and you should get k = w/r unless i am doing something wrong. r is the total radius of earth

OpenStudy (perl):

also there it is posted here https://www.physicsforums.com/threads/center-of-earth-work-problem-some-calculus.799459/

OpenStudy (perl):